MILLIMETER-WAVE SPECTROSCOPY OF TRANSIENT SPECIES IN SUPERSONIC JET

\underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

Submillimeter-wave absorption spectroscopy has been applied to observe transient species produced in the supersonic jet expansion with the UV laser photolysis, such as vinyl ^{a}, FeCO ^{b}, CoCO, CoNO, and FeNO.

For vinyl, the proton tunneling-rotation transitions, as well as the pure rotational transitions, have been observed for the H_{2}CCH, H_{2}CCD and D_{2}CCD isotopic species to determine tunneling splitting ΔE_{0} and the potential barrier height h for the proton (H/D) tunneling motion. Although these isotopic species have very different values for the tunneling splitting ΔE_{0} (16.185, 1.187 and 0.770 GHz), they have almost the same barrier heights h of 1580, 1520, and 1549 cm^{−1}. As for the HDCCH species, the pure rotational transitions (both a- and b-types) of only one isomer (cis-HDCCH:D cis to the unpaired electron) were observed in the jet cooled condition, but no lines for the trans-form nor cis-trans tunneling transitions were detected suggesting HDCCH has cis- and trans-isomers. The difference in zero point energy between the cis- and trans-HDCCH isomers is calculated to be 32 cm^{−1} with the CCSD(T)/aug-cc-pVTZ level ab initio calculation to support the present result. We also observed the large off-diagonal (ΔI = ±1) hyperfine interaction constants for H_{2}CCD which causes the nuclear spin conversion between the ortho (I_{β}=1) and para (I_{β}=0) H_{2}CCD.

The jet cooled FeCO, CoCO, FeNO, CoNO, and Co(CO)(NO) radicals were produced in a supersonic jet expansion by excimer laser photolysis of Fe(CO)_{5}, Co(CO)_{3}NO and Fe(CO)_{2}(NO)_{2} to observe the rotational lines. The molecular structures of FeCO, CoCO, CoNO and FeNO, for example, were confirmed to be linear with their electronic ground states of ^{3}Σ^{−}, ^{2}Δ_{i}, ^{1}Σ^{+}, and ^{2}Δ_{i} as suggested by high level ab initio calculations.

The setup was also used to observe the internal rotation transitions of the He-HCN ^{c}, H_{2}-HCN, Ne-HCN, Ar-HCN, and H_{2}-H_{2}O complexes in the SMMW range between 60 and 360 GHz. The potential energy surfaces calculated by CCSD(T) ab initio calculation were improved to explain the experimental results to the microwave accuracy.

aK. Tanaka, M. Toshimitsu, K. Harada, T. Tanaka, J. Chem. Phys. \underline{120} 3604, (2004).

bK. Tanaka, M. Shirasaka, T. Tanaka, J. Chem. Phys. \underline{106} 6820, (1997).

cK. Harada, K. Tanaka, T. Tanaka, S. Nanbu, M. Aoyagi, J. Chem. Phys. \underline{117} 7041, (2002).